what is the empirical formula of a compound that is 63.2% Mn and 36.8% O?
1
Expert's answer
2020-04-02T08:57:21-0400

Examle 1

For example, the mass of a substance = 100g, then m(Mn)=63.2g and m(O)=36.8g.


n(Mn)=m(Mn)/M(Mn)=63.2/55=1.15mol

n(O)=m(O)/M(O)=36.8/16=2.3mol


n(Mn)/n(O)=1.15/2.3=1/2


Empirical formula MnO2


Examle 2

w(Mn)=63.2% M(Mn)=55

63.2/55=1.15


w(O)=36.8% M(O)=16

36.8/16=2.3


1.15/1.15=1

2.3/1.15=2

Empirical formula MnO2